OpenStudy (anonymous):

list the coordinates of any two points on the graph in (x,y) form. the numbers on the horizontal axis are 0.5, 1.0, 1.5, 2.0, 2.5, 3.0, 3.5. 4.0. I have (2.0, 108000), (4.0, 96000). How to solve slope of this line

OpenStudy (anonymous):

slope is change in y over change in x. from 2 to 4 is two steps right. from 108000 to 96000 is 12000 steps down. slope is -6000 \[\frac{y_2-y_1}{x_2-x_1}=\frac{96000-108000}{4-2}=\frac{-12000}{2}=-6000\]